# LoaQ: Layer-wise Output Approximation Quantization

### v1

@@ Abstract
A natural and intuitive idea in model quantization is to approximate each component’s quantized output to match its original. Layer-wise post-training quantization (PTQ), though based on this idea, adopts a strictly local view and can achieve, at best, only activation-aware approximations of weights. As a result, it often leads to insufficient approximations and practical deviations from this guiding intuition. Recent work has achieved a more accurate approximation of linear-layer outputs within the framework of layer-wise PTQ, but such refinements remain inadequate for achieving alignment with the full model output. Based on a deeper understanding of the structural characteristics of mainstream LLMs, we propose LoaQ, an output-approximation method for layer-wise PTQ that explicitly targets output-level consistency. It better aligns with this intuition and can feature a simple closed-form solution, making it orthogonal to existing techniques and readily integrable into existing quantization pipelines. Experiments on the LLaMA and Qwen model families demonstrate that LoaQ performs effectively in both weight-only and weight-activation joint quantization. By integrating seamlessly with existing quantization strategies, it further enhances overall quantization quality and shows strong potential to advance the frontier of post-training quantization.
@@ 1 Introduction
In recent years, large-scale models have achieved remarkable breakthroughs in performance, yet these advances come with significant memory overhead, posing challenges for practical deployment. Model compression techniques (Zhu et al. 2024) have emerged as a promising solution to this problem, with post-training quantization (PTQ) (Krishnamoorthi 2018) being one of the most impactful approaches. Within PTQ methods, layer-wise PTQ has attracted particular attention. By formulating quantization at the level of individual linear layers, layer-wise PTQ simplifies both the process and the associated optimization objectives. Despite this simplification, layer-wise PTQ consistently delivers strong empirical performance. Moreover, this approach offers significant efficiency advantages: it can quantize large-scale models within a short time using only a single GPU. Beyond its standalone effectiveness, layer-wise PTQ also serves as an excellent initialization for end-to-end PTQ frameworks, accelerating their convergence and thereby further enhancing the final model’s performance. These properties collectively make layer-wise PTQ an important contributor to advancing the performance boundaries of quantized models.
@@
Although layer-wise PTQ methods, ranging from the classic GPTQ (Frantar et al. 2023) to the more recent MagR (Zhang et al. 2024), have achieved notable success, they still predominantly focus on weight approximation. This focus exposes a fundamental limitation: to maintain their layer-wise nature, these methods typically depend only on the input to each linear layer. Consequently, they are confined to activation-aware weight approximation, which diverges from the natural objective of output approximation. Recent approaches, including GPTAQ (Li et al. 2025), Qronos (Zhang et al. 2025), and QEP (Arai and Ichikawa 2025), overcome this limitation through an error-compensation strategy. Specifically, they modify weights to compensate for the mismatch between pre- and post-quantization inputs, while still following the forward layer-wise quantization sequence. By explicitly incorporating this compensation, these methods enable a more output-oriented optimization at the level of individual linear layers and ultimately achieve more accurate linear-layer output approximation.
@@
While previous methods approximate outputs at the linear-layer level, this scope remains confined and fails to account for residual connections and RMSNorm, making it insufficient for model-level output approximation. To address these issues, we extend our perspective to the sub-layer level, allowing us to naturally account for these components. Building on this idea, we introduce LoaQ, a layer-wise PTQ method that explicitly targets sub-layer output approximation, thereby overcoming the limitations of linear-layer methods and aligning more closely with model-level output approximation. Extensive experiments on LLaMA (Touvron et al. 2023; Grattafiori et al. 2024) and Qwen (Yang et al. 2025) model familes confirm that LoaQ consistently preserves high quantization accuracy across diverse tasks and quantization settings. Furthermore, LoaQ integrates seamlessly with other quantization strategies, including the Hadamard transform and NeUQI (Lin, Hu, and Wan 2025), and delivers additional improvements in overall model performance. Code and models will be released.
@@ 2 Preliminaries
2.1 Notation
In this section, we introduce the notation used consistently throughout this paper.
Let $W,Q\in\mathbb{R}^{N\times M}$ denote the original and quantized weight matrices, respectively. Similarly, let $X\in\mathbb{R}^{T\times N}$ and $\hat{X}\in\mathbb{R}^{T\times N}$ represent the activation matrices before and after quantization, where variables with a hat (e.g., $\hat{X}$) indicate quantized versions of their unhatted counterparts (e.g., $X$). Here, $T$ is the number of tokens, and $N$ and $M$ are the input and output feature dimensions, respectively.
@@ 2.1 Notation
For a vector $x\in\mathbb{R}^{N}$, define the normalized vector as
$\operatorname{Norm}(x)=r(x)\cdot x,$ (1)
where $r(x)=(N^{-1}\textstyle\sum_{i}x_{i}^{2}+\epsilon)^{-1/2}$, with $\epsilon>0$ added for numerical stability.
For a matrix $X\in\mathbb{R}^{T\times N}$, define the row-wise normalized matrix as
$\operatorname{Norm}(X)=R(X)X,$ (2)
where $R(X)=\operatorname{diag}\big{(}r(X_{1,:}),\dots,r(X_{T,:})\big{)}$.
2.2 Layer-wise PTQ
In model quantization, a common assumption is that accurately approximating the weights of each linear layer is sufficient to maintain overall model performance. Layer-wise PTQ builds on this principle by treating the quantization of each linear layer as an independent optimization task. To enhance efficiency, it employs a single-pass strategy that sequentially quantizes layers during one forward pass. Under this design, the GPU stores solely the minimal data necessary for the active layer, typically retaining only the post-quantization input for the optimization process. This approach substantially reduces computational overhead and memory consumption, resulting in a streamlined and efficient layer-wise PTQ framework.
In the weight-only quantization setting of layer-wise PTQ, most classical weight approximation objectives fall into two main forms.
  1. 1.
    Direct weight approximation minimizes
    $\arg\min_{Q}\|Q-W\|_{F}^{2},$ (3)
    where the objective seeks the optimal quantized value for each weight individually. A representative method under this formulation is Round-To-Nearest (RTN), which quantizes each weight by rounding it to the nearest value in the quantization grid.
  2. 2.
    Activation-aware weight approximation minimizes
    $\arg\min_{Q}\|\hat{X}(Q-W)\|_{F}^{2},$ (4)
    where the objective explicitly incorporates the input activation distribution, thereby mitigating the impact of quantization error on the linear layer output. A typical approach in this category is GPTQ, which leverages second-order information of this approximation objective to further reduce output distortion during quantization.
In joint weight and activation quantization, researchers have explored several strategies that vary in complexity and optimization scope: (i) independently quantizing weights and activations, following the principle of direct weight approximation; (ii) quantizing activations first, followed by activation-aware weight approximation to optimize the weights; (iii) jointly optimizing both weight and activation quantization to directly minimize the overall output error, which is often carried out based on the results of (i) or (ii).
Layer-wise PTQ can be further enhanced by integrating various complementary techniques. Notable examples include SmoothQuant (Xiao et al. 2023), which balances the numerical scales of weights and activations to alleviate quantization error; Hadamard transform, used in QuaRot (Ashkboos et al. 2024) and QuIP# (Chee et al. 2023), applies a Hadamard transformation to smooth the weight distribution and reduce quantization error; NeUQI (Lin, Hu, and Wan 2025), which improves the initialization of uniform quantization parameters; and BCD (Nair and Suggala 2024), which employs coordinate descent to refine the quantized weight of each linear layer.
Figure 1: The overall diagram of LoaQ. Each colored circle denotes an intermediate result, which LoaQ aims to approximate when quantizing the weight matrix of the corresponding color.The red circle 1 and green circle 2 correspond to the parts approximated by Linear-layer Output Approximation, the blue circle 3 corresponds to Sub-layer Output Approximation, and the orange circle 4 corresponds to Normalized Sub-layer Output Approximation.
2.3 Sub-layer
@@
In mainstream LLM architectures, each Transformer layer consists of a self-attention sub-layer and an MLP sub-layer. As shown in Figure 1, each sub-layer typically includes an RMSNorm, a sub-layer module, and a residual connection adding the input to the output. The sub-layer module can be uniformly represented in a simplified form:
$X_{k}^{\mathrm{out}}$ $=\phi\big{(}\mathrm{Norm}(h_{k})W_{\mathrm{in}}\big{)},$ (5)
@@
In this formulation:
  • For the self-attention module, $W_{\mathrm{in}}=[W_{q};W_{k};W_{v}]$ are the input projections to queries, keys, and values. The nonlinearity $\phi$ denotes the masked attention mechanism, which typically includes QK-normalization (if applicable), rotary positional encoding (RoPE), and causal attention over the resulting representations. The output projection is given by $W_{\mathrm{out}}=W_{o}$.
  • For the MLP module, $W_{\mathrm{in}}=[W_{\mathrm{gate}};W_{\mathrm{up}}]$ corresponds to the projections used in the gated feedforward path. The nonlinearity $\phi$ denotes an elementwise activation that gates the $W_{\mathrm{up}}$ projection using $W_{\mathrm{gate}}$ in the expanded hidden space. The projection back to the model dimension is given by $W_{\mathrm{out}}=W_{\mathrm{down}}$.
@@ 3 Methodology
In this section, we introduce LoaQ, short for Layer-wise Output Approximation Quantization. LoaQ builds on the layer-wise post-training quantization framework, which typically focuses on quantizing only $W_{\text{in}}$ and $W_{\text{out}}$. Extending this framework, LoaQ incorporates additional factors during quantization to achieve more accurate approximations. As shown in Figure 1, LoaQ consists of three components that progressively target the true output approximation with increasing accuracy: linear-layer, sub-layer, and sub-layer followed by RMSNorm. In contrast, GPTAQ and Qronos only target the linear-layer part. Each component in LoaQ builds upon the previous one, achieving increasingly accurate approximations by strategically selecting intermediate results within the model. The intermediate results highlighted by the red circle 1 and green circle 2 correspond to the linear-layer outputs, which are approximated using the method detailed in Section 3.1. With deeper insights into LLM architectures, the approach applied to the blue circle 3 replaces the approximation at the green circle 2 and is further enhanced by the method associated with the orange circle 4.
@@ 3.1 Linear Layer Output Approximation
In this subsection, we aim to make the results at the red circle 1 and the green circle 2 in Figure 1 after quantization closely match their pre-quantization counterparts by approximating the outputs of the corresponding linear layers. For each linear layer, this leads to the following optimization objective.
$\arg\min_{Q}\|\hat{X}Q-XW\|_{F}^{2}.$ (7)
Through a series of equivalent transformations, detailed in Appendix A, this is equivalent to:
$\arg\min_{Q}\|\hat{X}\big{(}Q-(I+H^{-1}C)W\big{)}\|_{F}^{2}.$ (8)
Therefore, updating $W$ to $(I+H^{-1}C)W$ adapts the problem to the activation-aware framework, allowing direct use of existing activation-aware weight approximation methods.
Figure 2: LLaMA 3.2 1B performance under different values of the adjustment parameter $\alpha$ with 3-bit channel-wise quantization.
Preliminary experiments on LLaMA 3.2 1B with 3-bit channel-wise quantization show that directly adding the linear-layer output approximation on top of the original GPTQ method leads to performance degradation, as demonstrated by the comparison between GPTQ and the Linear-Layer Output Approximation (LLOA) in Figure 2. Interestingly, for other methods based on linear-layer output approximations, such as GPTAQ and Qronos, we also observe similar degradation when they are applied solely to quantize the linear layers on top of the original layer-wise PTQ methods. To address Upon careful examination of the code, we found that these methods employ different strategies to mitigate the risk of large deviations. For instance, GPTAQ defines its objective as a weighted combination of the linear-layer output approximation objective and the original objective. Qronos, in contrast, resets the accumulated error at the start of each Transformer layer and focuses only on the newly introduced error within that layer. These strategies help ensure that the approximation target remains closely aligned with the true optimization objective.
Consequently, we adopt a more general technique by introducing a tunable adjustment parameter $\alpha\in[0,1]$, modifying the update rule to $W\leftarrow(I+\alpha H^{-1}C)W$. Figure 2, LLOA is indeed effective, but when the correction is too strong, it can deviate excessively from the final optimization objective.
QEP attributes this phenomenon to overfitting on the calibration set, as the number of calibration tokens is not much larger than the hidden state dimension. We instead identify a different cause. Although aligning linear-layer outputs with a surrogate objective is formally consistent with the final goal, residual errors can persist even under ideal conditions. Moreover, the surrogate objective may fail to capture the true problem structure, producing solutions that deviate from the real optimum and occasionally degrading performance under the final metric. These findings reveal the limitations of relying solely on such alignment constraints and naturally lead to the question of what in the LLM structure better represents the true optimization objective. Addressing this question is the focus of the following two sections.
3.2 Sub-layer Output Approximation
In this subsection, we shift the focus from approximating the green circle 2 to approximating the blue circle 3 in Figure 1, aiming to make the post-quantization results closely match their pre-quantization counterparts by aligning the outputs of the corresponding sub-layer.
Ideally, methods that approximate only the linear-layer output can at best match the output of the sub-layer module but fail to approximate the full sub-layer output $h_{k+1}$. This limitation arises because, by neglecting the residual connection, these methods overlook the accumulated errors propagated through the residual stream. In our approach, we aim to match the sub-layer output while preserving the layer-wise property. To this end, the sub-layer output approximation is only considered when quantizing $W_{\mathrm{out}}$. That is, we aim to approximate $h_{k+1}=h_{k}+X_{k}^{\mathrm{out}}W$ by $\hat{h}_{k}+\hat{X}_{k}^{\mathrm{out}}Q$. Accordingly, for notational simplicity, we omit both the subscript $k$ and the superscript $\mathrm{out}$ in the following formulation, and write the optimization objective as:
$\arg\min_{Q}\|(\hat{h}+\hat{X}Q)-(h+XW)\|_{F}^{2}$ (9)
@@ 3.2 Sub-layer Output Approximation
By transformation, we obtain the following equivalent form:
$\arg\min_{Q}\|\hat{X}\big{(}Q-\big{(}$ $(I+H^{-1}C)W+$
$H^{-1}\hat{X}^{\top}(h-\hat{h})\big{)}\big{)}\|_{F}^{2},$ (10)

### v2

@@ Abstract
A natural and intuitive idea in model quantization is to approximate each component’s quantized output to match its original. Motivated by this idea, most layer-wise post-training quantization (PTQ) methods focus on weight approximation at the linear-layer level. As a result, this local objective often yields insufficient approximations and practical deviations from the guiding intuition. Recent work has improved the approximation of linear-layer outputs within the layer-wise PTQ framework, but such refinements remain inadequate for achieving alignment with the full-model output. Based on a deeper understanding of the structure of mainstream LLMs, we propose LoaQ, which incorporates output-matching factors when quantizing linear layers within the layer-wise PTQ framework. It better aligns with this intuition and can feature a simple closed-form solution, making it orthogonal to existing techniques and readily integrable into existing quantization pipelines. Experiments on the LLaMA and Qwen model families demonstrate that LoaQ performs effectively in both weight-only and weight-activation quantization. By integrating seamlessly with existing quantization strategies, it further enhances overall quantization quality and shows strong potential to advance the frontier of post-training quantization. [^1] The code will be available for download
LoaQ: Layer-wise Output Approximation Quantization
Li Lin, Xiaojun Wan, Wangxuan Institute of Computer Technology, Peking University, efsotr_l@stu.pku.edu.cn, wanxiaojun@pku.edu.cn
@@ 1 Introduction
In recent years, large-scale models have achieved remarkable breakthroughs in performance, yet these advances come with significant memory overhead, posing challenges for practical deployment. Model compression techniques Zhu et al. (2024) have emerged as a promising solution to this problem, with post-training quantization (PTQ) Krishnamoorthi (2018) being one of the most impactful approaches. Within PTQ methods, layer-wise PTQ has attracted particular attention. By formulating quantization at the linear-layer level, layer-wise PTQ simplifies both the quantization process and the optimization objectives. Despite this simplification, layer-wise PTQ consistently delivers strong empirical performance. Moreover, this approach offers significant efficiency advantages: it can quantize large-scale models within a short time using only a single GPU. Beyond its standalone effectiveness, layer-wise PTQ also serves as an excellent initialization for end-to-end PTQ frameworks, accelerating their convergence and thereby further enhancing the performance. These properties collectively make layer-wise PTQ an important contributor to advancing the performance boundaries of quantized models.
@@
Although layer-wise PTQ methods, from classic GPTQ Frantar et al. (2023) to more recent MagR Zhang et al. (2024), have achieved notable success, they still primarily focus on weight approximation. As a result, to preserve the layer-wise setting, the only additional information they leverage is the input to each linear layer. Consequently, they are confined to activation-aware weight approximation, which diverges from the natural objective of output approximation. Recent approaches, including GPTAQ Li et al. (2025), Qronos Zhang et al. (2025), and QEP Arai and Ichikawa (2025), overcome this limitation through an error-compensation strategy. Specifically, they modify weights to compensate for the mismatch between pre- and post-quantization inputs. By explicitly incorporating this compensation, these methods enable a more output-oriented optimization at the linear-layer level and ultimately achieve more accurate linear-layer output approximation.
@@
While previous methods approximate outputs at the linear-layer level, this scope remains confined and fails to account for residual connections and RMSNorm, making it insufficient for model-level output approximation. To address these issues, we extend the perspective to the sub-block level, allowing us to naturally account for these components. Building on this idea, we introduce LoaQ, a layer-wise PTQ method that explicitly targets sub-block output approximation, thereby overcoming the limitations of linear-layer methods and aligning more closely with model-level output approximation. Extensive experiments on LLaMA Touvron et al. (2023); Grattafiori et al. (2024) and Qwen Yang et al. (2025) model families confirm that LoaQ consistently preserves higher quantization accuracy across diverse tasks and quantization settings. Furthermore, LoaQ integrates seamlessly with other quantization strategies, including the Hadamard transform and NeUQI Lin et al. (2025), and delivers additional improvements.
@@
2 Related Works
Layer-wise PTQ
Layer-wise PTQ independently optimizes each linear layer to preserve overall model performance, enabling simpler optimization procedures and more targeted algorithm design. Building on the layer-wise paradigm, prior work often improves PTQ by incorporating complementary techniques that reduce quantization error or stabilize optimization within each layer. Representative approaches include SmoothQuant (Xiao et al., 2023) and AWQ (Lin et al., 2024), which balance the numerical scales of weights and activations to alleviate quantization error; the Hadamard transform adopted in QuaRot (Ashkboos et al., 2024) and QuIP# (Chee et al., 2023) to smooth the weight distribution; NeUQI (Lin et al., 2025), which improves the initialization of uniform quantization parameters; and BCD (Nair and Suggala, 2024), which applies coordinate descent to refine the quantized weights of each linear layer.
Error-compensated PTQ
Methods such as GPTAQ Li et al. (2025), Qronos Zhang et al. (2025), and QEP Arai and Ichikawa (2025) address the fact that existing layer-wise PTQ methods fail to model the discrepancy between pre-quantization inputs and post-quantization inputs when quantizing a linear layer, resulting in accumulated error. However, these error compensation methods still do not consider the presence of residual connections, and thus, error accumulation along residual connection paths remains unresolved.
@@ 3 Preliminaries
3.1 Linear-layer Loss
The linear-layer loss is designed to guide the optimization toward reducing the quantization error in the linear layer output, formulated as
$\mathcal{L}(Q)=\|X(Q-W)\|_{F}^{2},$ (1)
where $W$ denotes the weight of the linear layer, $Q$ denotes its quantized counterpart, and $X$ denotes the input, also referred to as the activation, to the linear layer. Evidently, the linear-layer loss is independent across the columns of $W$, and thus can be decomposed as follows:
$\mathcal{L}(q)=\|X(q-w)\|_{2}^{2},$ (2)
where $w$ is a single column of $W$, and $q$ denotes its quantized counterpart. Moreover, $H=X^{\top}X$ is the Hessian matrix of the per-column linear-layer loss $\mathcal{L}(q)$.
Figure 1: The overall diagram of LoaQ. Each colored vertical bidirectional arrow denotes the corresponding approximate output obtained when quantizing the weight matrix of the same color.
3.2 GPTQ
To minimize the loss function in Eq. 1, the core idea of the GPTQ algorithm is to quantize the weight matrix sequentially in a row-wise manner, compensating for the quantization errors of previously processed rows using the remaining unquantized rows. The formulation is given as follows:
@@ 3.2 GPTQ
[…]
where $\mathcal{Q}$ denotes the quantization mapping, and $T$ denotes the compensation matrix, which is a strictly lower-triangular matrix that ensures each row accounts only for the quantization errors of the preceding rows. As stated in Chee et al. (2023), the optimal compensation matrix $T$ is $D^{-1}(L-D)$, where $L$ and $D$ are derived from the Cholesky decomposition of $H=LL^{\top}$ and $D$ is the diagonal matrix formed by the diagonal elements of $L$.
3.3 GPTAQ
Building on GPTQ, the GPTAQ algorithm (Li et al., 2025) extends this approach by introducing an compensation mechanism to better align quantized layer outputs with their original counterparts. The loss function for each linear layer is formulated as
$\mathcal{L}(Q)=\|XQ-X^{\prime}W\|_{F}^{2},$ (4)
where $X^{\prime}$ denotes the input to the original linear layer. The formulation of GPTAQ is
[…]
$\quad\quad+L(M_{L}\odot(L^{\top}(M_{L}\odot C)))\widetilde{W}$,
[…]
where $C=X^{\top}(X^{\prime}-X)$ serves as a correction term that compensates for the mismatch between the original and quantized layer inputs, $M_{L}$ denotes a strictly lower-triangular masking matrix with ones below the diagonal, and $\odot$ denotes element-wise multiplication. We can readily observe that the GPTAQ algorithm completely ignores the information in the upper-triangular part of $C$.
In addition, directly applying the formulation in Eq. 3.3 yields worse performance than GPTQ. Therefore, Li et al. (2025) introduce a tuning parameter $\alpha$, as shown in their GitHub repository.
[…]
$\quad\quad+\alpha L(M_{L}\odot(L^{\top}(M_{L}\odot C)))\widetilde{W}$,
[…]
3.4 Sub-block
@@
In mainstream LLM architectures, each Transformer block consists of a self-attention sub-block and an MLP sub-block. As shown in Figure 1, each sub-block typically includes an RMSNorm, a sub-block module, and a residual connection adding the input to the output. The sub-block can be uniformly represented in a simplified form:
$X_{k}^{\mathrm{out}}$ $=\phi\big(\mathrm{RMSNorm}(h_{k})W_{\mathrm{in}}\big),$ (7)
@@
For the self-attention module, $W_{\mathrm{in}}=[W_{q};W_{k};W_{v}]$ are the input projections to queries, keys, and values. The nonlinearity $\phi$ denotes the masked attention mechanism, which typically includes rotary positional encoding (RoPE), causal attention, and may also incorporate QK-Norm. The output projection is given by $W_{\mathrm{out}}=W_{o}$.
For the MLP module, $W_{\mathrm{in}}=[W_{\mathrm{gate}};W_{\mathrm{up}}]$ corresponds to the projections used in the gated feedforward path. The nonlinearity $\phi$ denotes an element-wise activation that gates the $W_{\mathrm{up}}$ projection using $W_{\mathrm{gate}}$ in the expanded hidden space. The projection back to the model dimension is given by $W_{\mathrm{out}}=W_{\mathrm{down}}$.
@@ 4 Methodology
In this section, we introduce LoaQ, short for Layer-wise Output Approximation Quantization. LoaQ builds on the layer-wise post-training quantization framework, which typically quantizes only $W_{\text{in}}$ and $W_{\text{out}}$, and extends it by incorporating additional factors during quantization to achieve more accurate approximations. As shown in Figure 1, LoaQ is organized into three components with hierarchical approximation objectives: linear-layer, sub-block, and sub-block followed by RMSNorm. The design transitions from linear-layer-level approximation to block-wise approximation, thereby imposing increasingly stringent constraints on aligning the quantized model with the original model.
@@ 4.1 Linear Layer Output Approximation
We aim to quantize the two linear layers $W_{\text{in}}$ and $W_{\text{out}}$ such that, as indicated by the red and green unidirectional arrows in Figure 1, the quantized outputs approximate their original counterparts as closely as possible. For each linear layer, we optimize the loss function in Eq. 4. By treating $X$ analogously to a common factor, we obtain an equivalent loss function for minimization:
$\mathcal{L}(Q)=\|X(Q-X^{\dagger}X^{\prime}W))\|_{F}^{2},$ (9)
where $X^{\dagger}=(X^{\top}X)^{-1}X^{\top}$ is left pseudoinverse. This loss function can be further simplified as
$\mathcal{L}(Q)=\|X\big(Q-(I+H^{-1}C)W\big)\|_{F}^{2}.$ (10)
Therefore, updating $W$ to $(I+H^{-1}C)W$ converts the problem into the loss function in Eq. 1, which can be directly solved using the GPTQ algorithm.
4.2 Sub-block Output Approximation
Ideally, methods that approximate only the linear-layer output can at best match the sub-block output. However, in the mainstream LLM architecture, the resulting errors accumulate through the residual connection. Accordingly, as indicated by the blue unidirectional arrow in Figure 1, LoaQ accounts for this effect when quantizing $W_{\mathrm{out}}$ by matching the sub-block output rather than only the linear-layer output. Given that $h$ and $h^{\prime}$ denote the hidden states of the quantized sub-block and the original sub-block, respectively, we obtain the loss function
$\mathcal{L}(Q)=\|(h+XQ)-(h^{\prime}+X^{\prime}W)\|_{F}^{2}.$ (11)
@@ 4.2 Sub-block Output Approximation
By applying the same procedure used to derive Eq. 10, we obtain the following equivalent form:
$\arg\min_{Q}\|X\big(Q-\big($ $(I+H^{-1}C)W+$
$H^{-1}X^{\top}(h^{\prime}-h)\big)\big)\|_{F}^{2}.$ (12)
Updating $W$ to $(I+H^{-1}C)W+H^{-1}X^{\top}(h^{\prime}-h)$ converts the problem into the loss function in Eq. 1.
Figure 2: Results under 3-bit quantization on LLaMA 3.2 1B, showing the mean squared error (MSE) of sub-blocks’ outputs. The left plot reports the MSE of unnormalized outputs, while the right plot reports the MSE of normalized outputs.
4.3 Normalized Sub-block Output Approximation
Based on the above understanding of the mainstream LLM architecture, the presence of residual connections leads to the accumulation of quantization error as the depth increases, as shown in the left plot of Figure 2. This raises the question of how the model remains stable despite the accumulation of quantization error. Upon closer examination, we observed that an RMSNorm layer is consistently applied before the input to each sub-block module and the final language model head. As illustrated in the right plot of Figure 2, this normalization mechanism prevents the mean squared error (MSE) between normalized outputs from growing beyond a certain depth.
Since a normalization layer follows each sub-block, it is the normalized output, rather than the raw sub-block output, that directly influences subsequent linear layers. Consequently, as indicated by the orange unidirectional arrow in Figure 1, aligning the normalized outputs of the quantized and original models becomes essential. To this end, we refine the objective to explicitly target the normalized sub-block outputs:
$\mathcal{L}(Q)=\|\rho(h+XQ)-\rho(h^{\prime}+X^{\prime}W)\|_{F}^{2},$ (13)
where $\rho(\cdot)$ is coloum-wise root mean square norm operator. To simplify the expression, we rewrite the normalization term as $\rho(h+XQ)=R(h+XQ)(h+XQ)$, where $R(\cdot)$ denotes the column-wise root mean square scaling operator. To decouple the quantized terms, we replace $Q$ with the original weight $W$ inside the scaling function $R(\cdot)$, and can then be transformed into the form:
$\mathcal{L}(Q)=\|\big(R(h+XW)h+(R(h+XW)X)Q\big)$
$-\big(R(h^{\prime}+X^{\prime}W)h+(R(h^{\prime}+X^{\prime}W)X^{\prime})W\big)\|_{F}^{2}$ (14)
This yields the same form as Eq. 11, allowing us to apply the same solution.
4.4 Tuning Parameters
Figure 3: Results under 3-bit quantization on LLaMA 3.2 1B, reporting average accuracy across different values of $\alpha$ and $\beta$.
For a single linear layer, the range of adjustable corrections is inherently limited. When the required correction is excessively large, it cannot be accommodated regardless of the adjustment strategy. In practice, the differences between hidden states are often large. When $h^{\prime}-h$ becomes too large, it is difficult to find a suitable $Q$ such that $h+XQ$ closely approximates $h^{\prime}+X^{\prime}W$. To address this issue, we introduce tuning parameters $\alpha$ and $\beta$ and change the update rule of $W$ to $(I+\alpha H^{-1}C)W+\beta H^{-1}X^{\top}(h^{\prime}-h)$. The detailed LoaQ pipeline is shown in Algorithm 1. As shown in Figure 3, the optimal performance is achieved when $\alpha$ takes intermediate values such as 0.4, 0.5, or 0.6, rather than 1.
Algorithm 1 LoaQ Pipeline
1:Weights $\{W_{\mathrm{in}}^{(\ell)},W_{\mathrm{out}}^{(\ell)}\}_{\ell=1}^{L}$, calibration inputs, bit‐width $k$, adjustment parameters $\alpha,\beta$
2:Quantized weights $\{Q^{(\ell)}\}_{\ell=1}^{L}$
3:Run forward to collect embedding outputs $h,h^{\prime}$
4:for $\ell=1$ to $L$ do
5:  for $W$ in $\{W_{\mathrm{in}}^{(\ell)},W_{\mathrm{out}}^{(\ell)}\}$ do
6:   Run forward to collect $X,X^{\prime}$
8:     $H\leftarrow X^{\top}X$
9:     $C\leftarrow X^{\top}(X^{\prime}-X)$ $\triangleright$ Correction term from Section 4.1
10:     $\Delta=\textbf{0}$